The length of a certain rectangle is twice the width. If the perimeter is 66 feet what is the length and the width?
Answer by checkley77(12844)

P=2L+2W
L=2W
66=2(2W)+2W
66=4W+2W
6W=66
W=66/6
W=11 FT. IS THE WIDTH.
L=2*11=22 FT. IS THE LENGTH.
PROOF:
66=2*22+2*11
66=44+22
66=66